OpenStudy (anonymous):

Asymphony has 125 musicians. Seventy five of them are women. Which of the following is the ration of the men to the total musicians in the symphony

OpenStudy (anonymous):

50:125

OpenStudy (stamp):

What an elegant explanation, UK_IN_US. I am sure that menure has a comprehensive understanding of ratios and their applications to problem solving. Bravo!

OpenStudy (stamp):

Ok but for real menure let's do this. Find the ratio of men to members. This is the problem, boiled down.

OpenStudy (stamp):

We know how many members there are (125) and how many women there are (75). To find the ratio of men to members\[men/members=\frac{men}{members}=men:members\] We need to know the number of men. We can substitute what we know: \[men/125=\frac{men}{125}=men:125\] But we do not yet know the number of men.

OpenStudy (anonymous):

I beg your pardon. Divide the number of men (125 (total) minus 75 ( women) = 25) by the total, 125.

OpenStudy (stamp):

Think logically, if men and women compose the orchestra, we can put that into math words by saying men + women = orchestra Again, let's substitute what we know. men + 75 = 125 This is a missing addend problem. Men is being added to 75 to make 125, so how many men are there?

OpenStudy (anonymous):

This gives you your ratio.

OpenStudy (stamp):

men + 75 = 125 men = 50 I hope you can figure out how I got 50 men. Either way, let's check. men + women = orchestra 50 + 75 = 125 125 = 125 It works So we have everything we need to finish our problem.

OpenStudy (stamp):

\[men/125=\frac{men}{125}=men:125\]men = 50 Let's put it in \[50/125=\frac{50}{125}=50:125\]

OpenStudy (anonymous):

Stamp, you explained it more comprehensively, I'll admit. However, the snide comment was not appreciated. Also, sorry to menure for turning this thread into a minor flaming.